OpenStudy (anonymous):

Please Help! Fan & medal :) What is the Axis of Symmetry Of Y=x^2+4x-2? & Y=-x^2+x-7?

jimthompson5910 (jim_thompson5910):

y = ax^2 + bx + c in general has the axis of symmetry to be x = -b/(2a)

jimthompson5910 (jim_thompson5910):

For example, say we had the equation y = 3x^2 - 6x + 10 a = 3 b = -6 c = 10 plug a = 3 and b = -6 into x = -b/(2a) to get x = -b/(2a) x = -(-6)/(2*3) x = 6/6 x = 1 So the axis of symmetry for the equation y = 3x^2 - 6x + 10 is x = 1

OpenStudy (ybarrap):

Couple more examples: